caffeine has a molecular weight 194.19g how much caffeine is required to make 1L of 0.1w/v caffeine


1
Expert's answer
2021-01-04T03:47:49-0500

Solution:


1) We should convert w/v to molarity:


"Molarity(\dfrac{mol}{L})=\dfrac{1000}{M_{solute}}\:\dfrac{\tfrac{w}{v}}{100}=\dfrac{1000}{194.19}\:\dfrac{0.1}{100}"


"C_M=0.00515\:(\tfrac{mole}{L})"

"0.00515=\dfrac{m_{caffeine}}{M_{caffeine}*V_{solution}}=\dfrac{m_{caffeine}}{194.19*1}"


"m_{caffeine}=1\:g"



Answer:

In order to make 1L of 0.1 "(\tfrac{w}{v})" caffeine, we need 1 g of it.